# Characterizing Density and Gravitational Potential Fluctuations of the Interstellar Medium

### v1

@@ 5.2 Comparison with Observations
Velocity-integrated 21 cm emission is the most direct tracer of gas surface density in regions that are dominated by atomic gas. For two-point statistics, power spectra of H I column densities have long been used to characterize the ISM, covering scales $10-10^{4}\,$pc in a range of nearby systems. For the SMC, S. Stanimirović et al. (1999) found an H I spectrum close to $P_{\delta}\propto k^{-3}$, somewhat steeper than our results. B. G. Elmegreen et al. (2001) found a steeper spectrum at small scales than large scales in the power spectrum of the LMC, suggesting that the transition marks scales comparable to the disk thickness. Based on analyses of several dwarfs and spirals, P. Dutta et al. (2009, 2013) also found a transition from a steeper spectrum at small scales to a shallower one at large scales. Making use of a large survey of dwarfs for the LITTLE THINGS project, measured (velocity-integrated) power spectra reported by H.-X. Zhang et al. (2012) were in the range $P_{\delta}\propto k^{-1.4}$ to $k^{-4.4}$ (without breaks), centered near $k^{-3}$, which is similar to (though a bit steeper than) our results. K. Grisdale et al. (2017) analyzed power spectra from H I in several massive disk galaxies, and found a range $P_{\delta}\propto k^{-1.6}$ to $k^{-2.8}$, with half having indices of $-2.1$ or $-2.2$ with no break, quite similar to our results. For M31 and M33, E. W. Koch et al. (2020) found indices $-2.7$ and $-2.1$, respectively, when fitting power spectra of H I maps at scales exceeding $\sim 200$ pc. K. Grisdale et al. (2017) and E. W. Koch et al. (2020) both conclude from detailed modeling that previous inferences of breaks in power spectra can be accounted for by the beam scale (or PSF) in observations. We do not observe a clear break in the power law index for our measured spectra, although there is a hint of a slightly shallower slopes at the largest scales we measure (approaching the box size of 1 kpc). In order to test definitively whether or not there is a break in the power spectrum (marking the gas scale height or other physical scale), TIGRESS simulations in boxes larger than those we analyze here would be required.
Dust emission also offers a promising means to analyze the detailed structure of the ISM, both from earlier multi-wavelength IR maps in individual nearby galaxies, and very recently with the advent of JWST surveying a large number of targets uniformly. E. W. Koch et al. (2020), analyzing power spectra of dust surface density maps in four Local Group galaxies (created by D. Utomo et al. 2019 from multi-wavelength Herschel data), concluded that a single power law with index between $-2$ and $-2.5$ describes the observations for the SMC, the LMC, and M31 well, while M33 has a shallower index. E. W. Koch et al. (2020) also measured power spectra in individual mid-to-far IR Herschel and Spitzer bands, finding indices similar to those for dust maps at long wavelength ($100-500\,\mu$m) and significantly shallower spectra at 24 $\mu$m, which they concluded was due to local heating in star-forming regions.
@@
Mid-IR emission maps from JWST for large spirals as well as dwarfs include both thermal dust-dominated and PAH-dominated bands, and go to very high spatial resolution. The analysis of 2D power spectra from 13 galaxies by C. Lind-Thomsen et al. (2025) produced a mean index $\sim-2$ in PAH-dominated bands, similar to the spectrum we find. Slopes measured in bands dominated by thermal dust were shallower ($\sim-1$ to $-1.5$), however. B. Elmegreen et al. (2025) found shallow slopes (index $\sim-1$) in both thermally dominated and PAH-dominated bands for a sample that included both spirals and dwarfs. While the latter work found no evidence for power spectrum breaks, the former did; since the sample overlaps, it is not clear what the origin of the discrepancy is.
@@
5.3 Characterizing Anisotropy and Non-Gaussianity
Given our model as specified in Section 3.3, a natural question is whether it fully encapsulates the structure of the fluctuation field. In particular, if a GRF with the specified power spectrum and random phases is generated, how do its properties differ from the simulated fluctuations? While the temporal phases of each $\mathbf{k}$-mode are reasonably well-described as uncorrelated (based on e.g., the lack of a nonzero peak frequency in Figure 8 and Figure 9), the spatial structure at a fixed time is more complex to describe. Figure 16 illustrates the difference between a snapshot of the actual logarithmic fluctuation $s$ in the R8 simulation (panel (a)), versus a snapshot as produced in a realization of a GRF with our model power spectrum using random phases for all Fourier components (panel (b)). While we see that the characteristic magnitudes (e.g., maxima and minima) of $s$ are similar across both fields, with similar overarching sizes of voids and overdensities, qualitative differences in the field structure are also evident. We attribute these differences to two approximations made in our models, which we describe below.
Figure 16: A comparison between a snapshot of the logarithmic surface density fluctuation $s$ for the TIGRESS R8 simulation (panel (a)) and a GRF generated using the model isotropic power law spectrum given in Equation 22 with parameters from Table 2 (panel (b)).
@@ 5.3 Characterizing Anisotropy and Non-Gaussianity
A second reason for the difference of the two panels in Figure 16 is that the model in Equation 22 treats the surface density fluctuation power as isotropic in $\mathbf{k}$-space. While this is true for intermediate to large $k$, for the smallest $k$ values that carry the most power, in reality the structures are primarily trailing (see Figure 6). This anisotropy also contributes to the larger scatter around the mean power spectrum present at the smallest $k$ values (see Figure 7). Nonetheless, isotropic models provide a reasonable description of the statistics, and are convenient in analytic applications due to their simplicity. In principle it is possible to characterize the degree of anisotropy or develop a precise theoretical model for its origin that specifies its characteristic tilt (see, e.g., T. Heinemann & J. C. B. Papaloizou 2009, 2012). We defer such an analysis and a characterization of the fluctuations that more accurately incorporates their slightly trailing nature to future work.
The importance of both the anisotropy and non-GRF-nature of ISM fluctuations is likely to vary depending on the application. For instance, current observational characterizations of power spectra in dust and gas emission, as discussed in Section 5.2, are presented as a function of the magnitude $k$ alone. In the stellar-dynamical context, the power spectrum of the potential entirely characterizes the dynamical influence of the fluctuations in quasilinear diffusion theory (J. Binney & C. Lacey, 1988; P.-H. Chavanis, 2023), and so the difference between the evolution of orbits under the real fluctuation field and a realization of a GRF with our model power spectrum may not be significant. We will assess the accuracy of this latter claim in a future work (Modak et al. in prep).
6 Summary
@@
    The pdf of the logarithmic surface density fluctuation $s\equiv\ln(\Sigma/\overline{\Sigma})$ is well approximated as a Gaussian (Equation 11 and Figure 3). When smoothed over larger scales, the variance is reduced, as expected given the power spectrum of the fluctuations (Section 3.1, Figure 5).
@@
    The spatial power spectrum of both $s$ and the linear surface density fluctuation $\delta\equiv(\Sigma-\overline{\Sigma})/\overline{\Sigma}$ are steep power laws, with fluctuation power concentrated at the largest scales (Figure 7, Equation 22, and Table 2). The power law indices we measure are similar to those measured in observed power spectra of gas and dust surface density fluctuations in external galaxies (Section 5.2).
@@
    The gravitational potential’s structure is well-described by a mean (independent of the in-plane coordinates) component that follows directly from the model for the vertical profile (Section 4.1, Figure 12), and a gravitational potential fluctuation power spectrum proportional to the surface density fluctuation spectrum (Equation 54, Figure 12, and Figure 13).
@@
    To generate a GRF realization for surface density fluctuations, Section 3.3 may be used (with $Q=s$ and parameters from Table 2 and Table 3). To generate a GRF realization of the corresponding volume density or gravitational potential, after specifying the time-dependence of $z_{\mathrm{com}}$ and $H$ (Figure 10, Table 4), the vertical profile model summarized in (iv) may be used in conjunction with Equation 49. Because spatial phases are correlated, it cannot be expected that maps of surface density generated in this way would resemble real maps. However, gravitational potential fluctuations generated in this way are all that is needed for applications to quasilinear evolution of the stellar distribution function.
@@
We thank Sanghyuk Moon, Chang-Goo Kim, Nickolas Kokron, Nora Linzer, Jacob Nibauer, David Thilker, Jenny Greene, and Jeremy Goodman for helpful conversations. S.M. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2039656. C.H. is supported by the John N. Bahcall Fellowship Fund and the Sivian Fund at the Institute for Advanced Study. Partial support for this work was provided by grant 510940 from the Simons Foundation to E.C.O.
@@ Appendix A Demonstrating Convergence with Simulation Resolution
In Figure 17 we plot the distribution of the logarithmic surface density fluctuation $p_{s}$ measured at each simulation’s respective grid scale for the R8 8 pc resolution (green) and 4 pc resolution (purple, as in the main text) simulations. As in Figure 3, the solid curve indicates the time-averaged quantity while the shaded region around it is a ±$\pm 1\sigma$ range around the average. We see that the distributions are both consistent with Gaussian distributions described by Equation 11 and Equation 12, with only small differences in the measured variances due to the different grid scales (which may be thought of as different smoothing scales for the same underlying field), as we discuss below.
@@
In Figure 18, analogous to Figure 5, we plot the standard deviation of the logarithmic surface density fluctuation field smoothed over a scale $R$, $\sigma_{s}(R)$, for a window function that is sharp in $k$-space (panel (a)) and sharp in real space (panel (b)), for the 4 pc resolution (purple) and 8 pc resolution (green) simulations. As expected for simulations converged in resolution, the standard deviation of the 4 pc simulation smoothed over an 8 pc scale is very close to that of the 8 pc simulation’s grid scale standard deviation (see the left-most portion of the $\sigma_{s}(R)$ curves in each panel). Indeed, across all smoothing scales, the curves lie nearly on top of each other, well within the ±$\pm 1\sigma$ shaded range. The variation around the time-averaged curve is larger for the lower resolution simulation despite using the same number of temporal snapshots, but this is to be expected since the total number of “samples” of the distribution is lower due to the smaller number of grid cells present.
@@
Finally, in Figure 19, analogous to Figure 7, we plot the power spectra $P_{Q}(k)$ for $Q=\delta$ (panel (a)) and $Q=s$ (panel (b)) for the 4 pc resolution (purple) and 8 pc resolution (green) simulations. At the largest scales (smallest $k$), the spectra coincide very well — we measure consistent normalizations and power law indices. Note that the reduced spatial resolution in the 8 pc simulation causes the spurious steepening of the spectrum to occur beginning at $kL\lesssim 100$ instead of $kL\approx 200$ for the 4 pc simulation, so to measure power law indices for that simulation we restrict to a smaller range than in the main text, $10\leq kL\leq 100$. To clearly identify the power law index over at least 1 dex, therefore, requires a simulation resolution of at least $\sim 8\,$pc, and we might expect the power law to extend to even higher $k$ in simulations with resolution finer than 444\,4pc. Regardless of whether the power law extends to smaller scales, the key conclusion that the power spectrum is featureless and most of the fluctuation power is present at the largest spatial scales continues to hold.
Appendix B Fitting a Fluctuation Vertical Profile

### v2

@@ V.2 Comparison with Observations
Velocity-integrated 21 cm emission is the most direct tracer of gas surface density in regions that are dominated by atomic gas. For two-point statistics, power spectra of H I column densities have long been used to characterize the ISM, covering scales $10-10^{4}\,$pc in a range of nearby systems. For the SMC, S. Stanimirović et al. (1999) found an H I spectrum close to $P_{\delta}\propto k^{-3}$, somewhat steeper than our results. B. G. Elmegreen et al. (2001) found a steeper spectrum at small scales than large scales in the power spectrum of the LMC, suggesting that the transition marks scales comparable to the disk thickness. Based on analyses of several dwarfs and spirals, P. Dutta et al. (2009, 2013) also found a transition from a steeper spectrum at small scales to a shallower one at large scales. Making use of a large survey of dwarfs for the LITTLE THINGS project, measured (velocity-integrated) power spectra reported by H.-X. Zhang et al. (2012) were in the range $P_{\delta}\propto k^{-1.4}$ to $k^{-4.4}$ (without breaks), centered near $k^{-3}$, which is similar to (though a bit steeper than) our results. K. Grisdale et al. (2017) analyzed power spectra from H I in several massive disk galaxies, and found a range $P_{\delta}\propto k^{-1.6}$ to $k^{-2.8}$, with half having indices of 2.1-2.1 or 2.2-2.2 with no break, quite similar to our results. For M31 and M33, E. W. Koch et al. (2020) found indices 2.7-2.7 and 2.1-2.1, respectively, when fitting power spectra of H I maps at scales exceeding $\sim 200$ pc.
While there have been some suggestions in the literature of breaks in the power spectrum appearing at scales corresponding to the disk thickness (i.e. $kH\sim\pi$), K. Grisdale et al. (2017) and E. W. Koch et al. (2020) both conclude from detailed modeling that previous inferences of breaks in power spectra can be accounted for by the beam scale (or PSF) in observations. We do not observe a clear break in the power law index for our measured spectra, although there may be a hint of a slightly shallower slope at the largest scales we measure (approaching the box size of $\sim 1$ kpc). In order to test definitively whether or not there is a break in the power spectrum (representing a transition in effective dimension for structures smaller than or larger than the gas scale height), TIGRESS simulations in boxes larger than those we analyze here would be required.
[…]
@@
Mid-IR emission maps from JWST for large spirals as well as dwarfs include both thermal dust-dominated and PAH-dominated bands, and go to very high spatial resolution. The analysis of 2D power spectra from 13 galaxies by C. Lind-Thomsen et al. (2025) produced a mean index $\sim-2$ in PAH-dominated bands, similar to the spectrum we find. Slopes measured in bands dominated by thermal dust were shallower ($\sim-1$ to 1.5-1.5), however. B. Elmegreen et al. (2025) found shallow slopes (index $\sim-1$) in both thermally dominated and PAH-dominated bands for a sample that included both spirals and dwarfs. While the latter work found no evidence for power spectrum breaks, the former did; since the sample overlaps, it is not clear what the origin of the discrepancy is.
@@
V.3 Characterizing Anisotropy and Non-Gaussianity
Given our model as specified in Section III.3, a natural question is whether it fully encapsulates the structure of the fluctuation field. In particular, if a GRF with the specified power spectrum and random phases is generated, how do its properties differ from the simulated fluctuations? While the temporal phases of each $\mathbf{k}$-mode are reasonably well-described as uncorrelated (based on e.g., the lack of a nonzero peak frequency in Figure 8 and Figure 9), the spatial structure at a fixed time is more complex to describe. Figure 16 illustrates the difference between a snapshot of the actual logarithmic fluctuation $s$ in the R8 simulation (panel (a)), versus a snapshot as produced in a realization of a GRF with our model power spectrum using random phases for all Fourier components as described in Section B.1 (panel (b)). While we see that the characteristic magnitudes (e.g., maxima and minima) of $s$ are similar across both fields, with similar overarching sizes of voids and overdensities, qualitative differences in the field structure are also evident. We attribute these differences to two approximations made in our models, which we describe below.
Figure 16: A comparison between a snapshot of the logarithmic surface density fluctuation $s$ for the TIGRESS R8 simulation (panel (a)) and a snapshot of a GRF generated using the model power spectrum following the method described in Section B.1 with parameters from Table 2 and Table 3 (panel (b)).
@@ V.3 Characterizing Anisotropy and Non-Gaussianity
A second reason for the difference of the two panels in Figure 16 is that the model in Equation 22 treats the surface density fluctuation power as isotropic in $\mathbf{k}$-space. While this is true for intermediate to large $k$, for the smallest $k$ values that carry the most power, in reality the structures are primarily trailing: ellipses fitted to the contours shown in Figure 6 monotonically decrease in eccentricity from $e\approx 0.9$ at the largest scales to $e\approx 0.2$ at the smallest scales. We note that the pitch angle (measured as the orientation of the major axis of the fitted ellipses) also varies with scale. These deviations from isotropy ($e=0$) contribute to the larger scatter around the mean power spectrum present at the smallest $k$ values (see Figure 7). Nonetheless, isotropic models provide a reasonable description of the statistics, and are convenient in analytic applications due to their simplicity.
Potentially, it may be possible to develop a theoretical model for the spectrum that specifies its characteristic anisotropy and tilt as a function of scale. Because the self-gravity of the gas is modest (the typical Toomre $Q$ parameter ranges from $\sim 1.2-3$ for the models studied here), mode amplitudes primarily reflect the magnetohydrodynamic response of the sheared system to turbulence excited by feedback, though self-gravitating swing amplification may still play a role (see, e.g., T. Heinemann & J. C. B. Papaloizou 2009, 2012). Given the strong nonlinearities and challenge of properly specifying excitation for a highly inhomogeneous system, however, development of rigorous theory is nontrivial. We defer such an analysis to future work.
The importance of both the anisotropy and non-GRF-nature of ISM fluctuations is likely to vary depending on the application. For instance, current observational characterizations of power spectra in dust and gas emission, as discussed in Section V.2, are presented as a function of the magnitude $k$ alone. In the stellar-dynamical context, the power spectrum of the potential entirely characterizes the dynamical influence of the fluctuations in quasilinear diffusion theory (J. Binney & C. Lacey, 1988; P.-H. Chavanis, 2023), and so the difference between the evolution of orbits under the real fluctuation field and a realization of a GRF with our model power spectrum may not be significant. We will assess the accuracy of this latter claim in a future work (Modak et al. in prep).
VI Summary
@@
    The pdf of the logarithmic surface density fluctuation $s\equiv\ln(\Sigma/\overline{\Sigma})$ is well approximated as a Gaussian (Equation 11 and Figure 3). When smoothed over larger scales, the variance is reduced, as expected given the power spectrum of the fluctuations (Section III.1, Figure 5).
@@
    The spatial power spectrum of both $s$ and the linear surface density fluctuation $\delta\equiv(\Sigma-\overline{\Sigma})/\overline{\Sigma}$ are steep power laws, with fluctuation power concentrated at the largest scales (Figure 7, Equation 22, and Table 2). The power law indices we measure are similar to those measured in observed power spectra of gas and dust surface density fluctuations in external galaxies (Section V.2).
@@
    The gravitational potential’s structure is well-described by a mean (independent of the in-plane coordinates) component that follows directly from the model for the vertical profile (Section IV.1, Figure 12), and a gravitational potential fluctuation power spectrum proportional to the surface density fluctuation spectrum (Equation 54, Figure 12, and Figure 13).
@@
    To generate a GRF realization for surface density fluctuations, as detailed in Appendix B, Section III.3 may be used (with $Q=s$ and parameters from Table 2 and Table 3). To generate a realization of the corresponding volume density or gravitational potential, after specifying the time-dependence of $z_{\mathrm{com}}$ and $H$ (Figure 10, Table 4), the vertical profile model summarized in (iv) may be used in conjunction with Equation 49. Because spatial phases are correlated, it cannot be expected that maps of surface density generated in this way would resemble real maps. However, gravitational potential fluctuations generated in this way are all that is needed for applications to quasilinear evolution of the stellar distribution function.
@@
We thank Sanghyuk Moon, Chang-Goo Kim, Nickolas Kokron, Nora Linzer, Jacob Nibauer, David Thilker, Jenny Greene, and Jeremy Goodman for helpful conversations. We also thank the referee for a thorough and insightful report. S.M. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2039656. C.H. is supported by the John N. Bahcall Fellowship Fund at the Institute for Advanced Study. Partial support for this work was provided by grant 510940 from the Simons Foundation to E.C.O.
@@ Appendix A Demonstrating Convergence with Simulation Resolution
In Figure 17 we plot the distribution of the logarithmic surface density fluctuation $p_{s}$ measured at each simulation’s respective grid scale for the R8 8 pc resolution (green) and 4 pc resolution (purple, as in the main text) simulations. As in Figure 3, the solid curve indicates the time-averaged quantity while the shaded region around it is a $\pm 1\sigma$ range around the average. We see that the distributions are both consistent with Gaussian distributions described by Equation 11 and Equation 12, with only small differences in the measured variances due to the different grid scales (which may be thought of as different smoothing scales for the same underlying field), as we discuss below.
@@
In Figure 18, analogous to Figure 5, we plot the standard deviation of the logarithmic surface density fluctuation field smoothed over a scale $R$, $\sigma_{s}(R)$, for a window function that is sharp in $k$-space (panel (a)) and sharp in real space (panel (b)), for the 4 pc resolution (purple) and 8 pc resolution (green) simulations. As expected for simulations converged in resolution, the standard deviation of the 4 pc simulation smoothed over an 8 pc scale is very close to that of the 8 pc simulation’s grid scale standard deviation (see the left-most portion of the $\sigma_{s}(R)$ curves in each panel). Indeed, across all smoothing scales, the curves lie nearly on top of each other, well within the $\pm 1\sigma$ shaded range. The variation around the time-averaged curve is larger for the lower resolution simulation despite using the same number of temporal snapshots, but this is to be expected since the total number of “samples” of the distribution is lower due to the smaller number of grid cells present.
@@
Finally, in Figure 19, analogous to Figure 7, we plot the power spectra $P_{Q}(k)$ for $Q=\delta$ (panel (a)) and $Q=s$ (panel (b)) for the 4 pc resolution (purple) and 8 pc resolution (green) simulations. At the largest scales (smallest $k$), the spectra coincide very well—we measure consistent normalizations and power law indices. Note that the reduced spatial resolution in the 8 pc simulation causes the spurious steepening of the spectrum to occur beginning at $kL\lesssim 100$ instead of $kL\approx 200$ for the 4 pc simulation, so to measure power law indices for that simulation we restrict to a smaller range than in the main text, $10\leq kL\leq 100$. To clearly identify the power law index over at least 1 dex, therefore, requires a simulation resolution of at least $\sim 8\,$pc, and we might expect the power law to extend to even higher $k$ in simulations with resolution finer than 44\,pc. Regardless of whether the power law extends to smaller scales, the key conclusion that the power spectrum is featureless and most of the fluctuation power is present at the largest spatial scales continues to hold.
Appendix B Generating Realizations of Surface Density Fluctuations
In this appendix, we describe two methods for generating realizations of an ISM surface density fluctuation field $Q(\mathbf{x},t)$ (where $Q=s$, $\delta$) using the model given in Section III.3 for the spatio-temporal power spectrum of the fluctuations. Both methods described here produce a fluctuation field with spatio-temporal Fourier amplitudes that match those of the simulated fluctuations (to the extent that the model matches the measured spectrum), but do not attempt to match the Fourier phases of the simulated fluctuation field. Since the fluctuations are highly nonlinear in amplitude, phases are correlated in a nontrivial way. As a result, model fluctuations generated by both methods lack many of the visual characteristics of the simulated ISM, like filamentary features (see also Section V.3 and Figure 16). However, these model fluctuations still prove useful for capturing the ISM-driven orbital heating/migration, as we show in Modak et al. (in prep.). This is what one expects from quasilinear diffusion theory, where phase information can be neglected (J. Binney & C. Lacey, 1988; C. Hamilton & J.-B. Fouvry, 2024).
B.1 Producing a Spatio-temporal Gaussian Random Field
The most direct method for generating a time series of fluctuation snapshots given their spatio-temporal power spectrum is to approximate the fluctuations as a GRF. A field produced in this way will have a Gaussian pdf, so it is most physically appropriate to use this method for a field $Q=s$. The standard algorithm for producing a realization of a GRF with power spectrum $P_{Q}(k,\omega)$ on a finite grid of size $N_{x}$, $N_{y}$, and $N_{t}$ in the $x$, $y$, and $t$ directions respectively is as follows:
  1. (i)
    Sample $N_{x}\cdot N_{y}\cdot N_{t}$ values from a standard normal distribution, and reshape the values to the desired shape of the full sequence of snapshots of the fluctuation field, i.e., assign spatial and temporal coordinates to each sample to produce a field $f(\mathbf{x},t)$ with a Gaussian pdf.
  2. (ii)
    Fourier transform the resulting field in both space and time to produce $f_{\mathbf{k}}(\omega)$.
  3. (iii)
    Multiply the field values $f_{\mathbf{k}}(\omega)$ by $\sqrt{P_{Q}(k,\omega)}$.
  4. (iv)
    Invert the Fourier transform to arrive at a field realization $Q(\mathbf{x},t)$ with the desired spectrum $P_{Q}$.
Implicit in the use of the Fourier transform in steps (ii) and (iv) is the assumption that the field satisfies periodic boundary conditions in space. However, the ISM fluctuations studied in this work instead satisfy shearing-periodic boundary conditions as given in Equation 3. Thus, the standard GRF realization algorithm must be modified as follows:
  1. (i)
    Follow steps (i)-(iii) as described above.
  2. (ii)
    Instead of performing the inverse Fourier transform in both space and time as described in (iv) above, first, only the temporal dimension should be inverse transformed, to produce $Q_{\mathbf{k}}(t)$.
  3. (iii)
    Convert from the unsheared wavevectors $\mathbf{k}$ to the sheared wavevectors $\mathbf{k}^{\prime}$ using the time-dependent relations given in Equation 4 and Equation 5.
  4. (iv)
    Inverse Fourier transform the resulting field to arrive at the field $Q(\mathbf{x}^{\prime},t)$ in sheared spatial coordinates.
  5. (v)
    Convert from the sheared coordinates $\mathbf{x}^{\prime}$ to the unsheared coordinates $\mathbf{x}$ using the time-dependent relations given in Equation 1 and Equation 2 to arrive at a realization $Q(\mathbf{x},t)$.
By separately performing the temporal and spatial transforms in this way, the spatio-temporal power spectrum remains unchanged, but the field now satisfies spatial shearing-periodic boundary conditions as desired. Note that the use of the time-dependent transformation from sheared to unsheared coordinates requires the choice of a temporal origin, and the snapshots will include a discontinuous jump in time-evolution at each “nearest periodic time” $t_{n}=nL_{y}/(q\Omega L_{x})$, because the shearing transform depends on $(t-t_{n})$ and $n=\lfloor q\Omega tL_{x}/L_{y}\rceil$ is a discontinuous function of time. However, in preliminary numerical tests of the dynamical influence of ISM fluctuations produced using this method, we find this point discontinuity to have no effect on the ensemble-averaged evolution of stellar orbits (Modak et al., in prep).
B.2 Spatial GRFs with Exponential Time-Dependence
The temporal discontinuity described above occurs because of the dependence on $t-t_{n}$, which is necessary to prevent fluctuations winding to very high azimuthal wavevectors $k_{y}$. However, this is simply a shortcoming of the GRF realization algorithm: in reality, fluctuation structures are disrupted by feedback processes on timescales much shorter than the winding time. An alternative method, which avoids these discontinuities, is to use the time-dependence specified directly in Equation 31. The procedure is as follows:
  1. (i)
    Choose $N_{p}$ “peak times” $\{t_{j}\}$ uniformly at random from the interval $t_{\mathrm{buffer}},t_{\mathrm{sim}}+t_{\mathrm{buffer}}]$ where $t_{\mathrm{sim}}$ is the desired temporal baseline for the fluctuations, and the extension of the interval by $t_{\mathrm{buffer}}$ on either side allows for modes to be rising or falling in amplitude even at the boundaries of the time interval of interest.
  2. (ii)
    For each peak time $t_{j}$, perform the following steps to generate a (purely spatial) GRF:
    1. (a)
      Sample $N_{x}\cdot N_{y}$ values from a standard normal distribution, and reshape the values to the desired shape of a single snapshot of the fluctuation field, i.e., assign spatial coordinates to each sample to produce an uncorrelated field $f^{(j)}(\mathbf{x})$.
    2. (b)
      Fourier transform the resulting field in space to produce $f^{(j)}_{\mathbf{k}}$.
    3. (c)
      Multiply the field values $f_{\mathbf{k}}$ by $\sqrt{P_{Q}(k)\omega_{0}(k)t_{\mathrm{sim}}/(4N_{p})}$ to produce a realization of a single structure $Q^{(j)}_{\mathbf{k}}$. In addition to the spatial power spectrum, the factor of $\omega_{0}(k)t_{\mathrm{sim}}/(4N_{p})$ ensures the resulting superposition of structures is appropriately normalized.
  3. (iii)
    For each desired snapshot time $t$, combine the individual GRFs $Q^{(j)}_{\mathbf{k}}$ as follows:
    1. (a)
      Multiply the values $Q^{(j)}_{\mathbf{k}}$ by $e^{-\omega_{0}(k)|t-t_{j}|}$ to produce each structure’s time-dependent contribution to the full fluctuation field, $Q^{(j)}_{\mathbf{k}}(t)$.
    2. (b)
      Convert from the unsheared wavevectors $\mathbf{k}$ to the sheared wavevectors $\mathbf{k}^{\prime}$ using the time-dependent relations given in Equation 4 and Equation 5, except using $t$ in place of $(t-t_{n})$, e.g., $k_{x}^{\prime}=k_{x}-q\Omega tk_{y}$, to produce each structure’s contribution in sheared wavevectors, $Q^{(j)}_{\mathbf{k}^{\prime}}(t)$.
    3. (c)
      Sum the contributions to arrive at $Q_{\mathbf{k}^{\prime}}(t)\equiv\sum_{j}Q^{(j)}_{\mathbf{k}^{\prime}}(t)$.
  4. (iv)
    Inverse Fourier transform the resulting field to arrive at the field $Q(\mathbf{x}^{\prime},t)$ in sheared spatial coordinates.
  5. (v)
    Convert from the sheared coordinates $\mathbf{x}^{\prime}$ to the unsheared coordinates $\mathbf{x}$ using the time-dependent relations given in Equation 1 and Equation 2, except using $t$ in place of $(t-t_{n})$ as in (iiib) above, to arrive at a realization $Q(\mathbf{x},t)$.
This method correctly reproduces the model spatio-temporal power spectrum and does not include any temporal discontinuities. The price we pay is that we have introduced two new free parameters $N_{p}$ and $t_{\mathrm{buffer}}$. Empirically, we find that setting $N_{p}\sim t_{\mathrm{sim}}/\tau_{0}$ and $t_{\mathrm{buffer}}\sim 2\tau_{0}$ yields fluctuation fields with spatio-temporal power spectra that are well-converged to the values measured in the TIGRESS simulations; these choices ensure that on average, at least one large-scale structure is near its peak amplitude for each simulation snapshot (see Equation 27). We note that in preliminary numerical tests, we find no significant differences between the dynamical influences of ISM fluctuations produced using this method and the GRF method described in the previous subsection (Modak et al., in prep).
Appendix C Fitting a Fluctuation Vertical Profile